The annual precipitation amounts in a certain mountain range are normally distributed with a mean of 88 inches, and a standard deviation of 10 inches. What is the likelihood that the mean annual precipitation during 25 randomly picked years will be less than 90.8 inches?
Answer by stanbon(75887)

The annual precipitation amounts in a certain mountain range are normally distributed with a mean of 88 inches, and a standard deviation of 10 inches. What is the likelihood that the mean annual precipitation during 25 randomly picked years will be less than 90.8 inches?
-----------------
z(90.8) = (90.8-88)/[10/sqrt(25)] = 2.8/2 = 1.4
---------------------------
P(x < 90.8) = P(z < 1.4) = normalcdf(-100,1.4) = 0.9192
===================
Cheers,
Stan H.
-----------------